# Equilibrium in Style: A Modeling Framework on the Cash Flow and the Life Cycle of a Consumer Store

### v1

@@ Abstract
The consumer store is ubiquitous and plays an important role in our everyday lives. It is an open question why stores usually have such short life cycles (typically around 3 years in China). This paper proposes a theoretical framework based on an equilibrium in style supply of stores and style demand of consumers to characterize store cash flow (revenue), leading to a strong explanation of this puzzle. In our model, we derive that the preference shifting of consumers is the main reason for the cash flow decreasing to its break-even line over time, while the visibility broadening leads to initial growth, resulting in rainbow-shaped cash flow and its life cycle. Moreover, the intensified spatial competition will lead to an unexpected decrease in the store’s cash flow, or even closure. We calibrate our model with proprietary data of three Chinese stores from three representative industries and study the relationship between customers’ preference shifting and cash flow. To our knowledge, there have been no prior attempts to quantitatively model the life cycle of the store.
@@ 1 Introduction
In the vast forest of the market place, conventional listed companies are the towering trees, everlasting and long-living, while consumer stores (micro-enterprises) embody the plentiful small plants which naturally grow and die with the changing seasons. These stores, such as convenience stores, restaurants, and hotels, exist everywhere in our daily life, little is known about them academically. Especially their frequent opening and closing, leaving a question of why these stores always have short lifespan. This paper develops a framework based on equilibrium in style to model stores’ cash flow over time, which can be calibrated by real data, and hence provides a capable explanation for this puzzle.
@@
The average level of cash flow determines the life cycle of a store as it faces some necessary costs, such as costs of raw materials, rents, and salary [^1] In this paper, we do not distinguish the two words: the cash flow and the revenue.. Usually one who passes by the store has a certain probability to walk into the store and consume goods and services (thereby the person is converted into an effective consumer). From the theoretical perspective, our model characterizes the cash flow of one store via the following three steps, among which we propose novel concepts and assumptions to precisely characterize our framework.
@@
First, we give a mathematical definition of cash flow and break down it forming the product of the average consumer price (ACP), the conversion rate, and the potential customer flow. The conversion rate is an important topic in economics. Perdikaki et al. (2012) define the conversion rate as the ratio of the number of transactions to that of the traffic and study the impact of traffic on cash flow. Furthermore, to better understand the conversion rate in our paper, we define the concept of population structure: classifying the population into several preference groups based on their age, occupation, and income, with each group of the population having its own conversion rate. Moreover, we weigh the overall conversion rate by the conversion rate of each group.
@@
Second, the preference is defined with two things in mind: the demand of the consumer and the supply of the store. On one hand, we capture the preference shifting of consumers by using the utility function in Cobb and Douglas (1928) and the Berry-Levinsohn-Pakes (BLP) model in Berry et al. (1995). The latter model was the first to propose the utility of different types of consumers concerning the product style. The model of the time-varying form is later taken into account in Berry and Haile (2014). On the other hand, we further define the product style strategy of the store. We assume that the product style and ACP of a given store are stable over time due to technical stickiness. As a result, we find a Nash equilibrium, where we derive the mathematical form of the conversion rate for each group of people.
Third, we turn to model the flow of potential customers. We assume the population is uniformly distributed around the store location with a constant population flow density. Furthermore, similar to Lucas and Rossi-Hansberg (2002), we assume the store’s attractiveness exponentially decreases with distance increasing. Additionally, at the beginning of a store opening, we assume no one knows the store and its visibility is broadening by a certain speed.
@@
Finally, we obtain an equation of the store’s cash flow over time. In particular, we derive that when consumer preferences are not dispersed, the cash flow of the store over time is a rainbow-shaped curve. The increase in the cash flow curve can be interpreted as increases of potential customers due to the broadening of the store’s visibility, while the decrease can be interpreted as decreases in the conversion rate due to a shift in consumer preferences. As a conclusion, we provide a road map for our modeling framework in Figure 1.
@@
Cash flow $\text{CF}_{t}$ Average consumer price $\theta_{t}$ Preferences Potential consumer flow $N_{t}$ Product price $\boldsymbol{\theta}^{*}(\boldsymbol{z})$ Population flow density $u$ Attraction $\boldsymbol{q}_{t}$ Store competition Visibility broadening $k$ Attractiveness decrease $e^{-\delta\|\mathbf{x}\|}$ Consumer population structure $\mathbf{P}$ Score $F$, Utility $U$ Store style Initial investment $I$, style set $\mathbb{Z}_{t}$ Equilibrium Purchase probability $\rho^{*}$ Product and storefront attributes $\mathbf{z}^{*}=(\mathbf{x},\boldsymbol{\xi})$ Life cycle $T$ Consumer conversion rate $\tilde{\beta}_{t}$ Preference shifting $\nu$
@@
In the empirical study, we calibrate our model by simulation method to illustrate some important implications and use the real data from three Chinese stores from three industries in the real economy: the retail industry, the restaurant industry, and the service industry. The results show that our model provides a strong explanation for the puzzle. First, the simulation results indicate that (i) the visibility broadening speed mainly affects the store’s upfront cash flow; (ii) the decrease coefficient mainly affects the length of the store’s life cycle; and (iii) the population flow density mainly affects the macro-scale of cash flow over time. Second, by using the real data of three Chinese stores, we conduct an Non-linear Least Squares regression (NLS) of the adjusted cash flow on the time scale to evaluate the the visibility broadening speed, the initial conversion rate and the decrease coefficient. The regression results uncover that the estimates are all statistically significant at a 1% confidence level. Finally, we plot the real data and the fitted data to illustrates that our model well captures the trend of the store’s cash flow, which strongly justifies the explanation of our model.
Furthermore, intensive spatial competition leads to an unexpected drop in the store’s cash flow. To better illustrate, the debut of a new same-brand store near the initial store often prompts consumer defection. In cases where the local demand cannot sustain both establishments, the closure of one of the stores becomes inevitable.
Our model incorporates several key ingredients of interest in literature, including product specialization, customers’ preferences, and purchase conversion rate. Burdett and Judd (1993) establish the equilibrium for dispersed prices under certain cases. Menzio (2023) discusses the relationship between product specialization to customers’ preferences, search friction decrease, and price dispersion. Menzio (2023) finds that the increase in product specialization helps sellers cater towards customers’ preference shifting, which leads to a higher store profit and also a higher customer surplus. The search friction can be seen as the visibility broadening in our paper. Kashyap (1995) uses real data to illustrate the effects of the size, frequency, and synchronization of price changes for twelve selected retail goods. In the field of marketing science, Lam et al. (2001) propose a framework to carefully classify the conversion rate into three effects of attraction (store-entry), conversion (purchase decision), and spending (purchase amount). They formulate a joint model of four simultaneous equations using front traffic, store traffic, number of store transactions, and store sales as the endogenous variables. Zimmermann and Auinger (2023) study a digital retailer and adopt a conversion rate optimization framework to increase sales.
The paper is organized as follows. Section 2 proposes the model environment of our paper. The demand of consumers and the supply of store are well defined in Section 3. Section 4 defines the concept of preference shifting and derive a Nash equilibrium. The spatial competition is described and we close the model in Section 5. The full model is presented in Section 6. We provide empirical evidence in Section 7. Section 8 provide the further discussions. Section 9 concludes.
@@ 2 Model Environments
Store. We study the typical consumer store in this paper. A (consumer) store is a type of micro-enterprise, usually possessing the following features: (i) Initial monetary investment is small. For example, it costs the owner around 40,000 Chinese yuan to invest in a luckin coffee store in Beijing, China. (ii) Routine expenditures include rent and salary, which forms a surprisingly stable ratio to daily income. For instance, the ratio of daily rent to daily cash flow is around 6.17% for a typical Chinese convenience store on 2022, according to the annual report of KPMG and China Chain-Store & Franchise Association (CCFA). (iii) The store has a high-frequency cash flow. For example, a Family Mart store typically has 1,000 orders (purchases) per day. (iv) The store caters to people’s daily demand, which encompasses the restaurant, retail, and service industry. (v) Most consumers of stores purchase goods and service in-store, especially the stores in the service industry. Therefore, the location of the store impacts their financial performance. (vi) There exists perfect competition (over 100 million stores in China) in each industry. In short, our focus is not on the conventional company, but the store, which are usually the following three types: direct-sale store, franchised outlet, and “mom-and-pop store”.
Cash flow decomposition. In this paper, we fix one store on time set $\mathcal{T}.$ The cash flow CF is a non-negative function on $\mathcal{T}.$ Then, we model the cash flow at time $t$ by a product of the average customer price (ACP) $\theta_{t}$, flow of potential customers $N_{t}$ and the conversion rate $\beta_{t}$, which can be formulated as
@@
Attributes and Style. Consider an economy corresponding to a consumer industry on time set $\mathcal{T}=\mathbb{R}.$ At any time $t\in\mathcal{T}$, there are new stores opened with different initial investments. Let the set of all possible initial investments be $\mathcal{I}=[I_{m},I_{M}]\subset\mathbb{R}^{+},$ which is a bounded set of a non-zero measure that does not change over time. Therefore, stores are infinitely numerous in this economy. The style of a store is the combination of its products and storefront. We assume that a store provides one representative product (commodity), characterized by a constant price $\theta$ and product attributes $\boldsymbol{x}\in\mathbb{R}^{p}$, which reflect the features of product size, taste, shape, etc. Similar assumption is mentioned or implied in Chamberlin (1933), Dixit and Stiglitz (1977) and Romer (1986). Moreover, we denote the storefront attributes by $\boldsymbol{\xi}\in\mathbb{R}^{q}$, which reflect the implicit influencing factors of store service, decoration, reputation, etc. We define the style vector for a store $\boldsymbol{z}=(\boldsymbol{x}^{\prime},\boldsymbol{\xi}^{\prime})^{\prime},$ and $\mathbb{R}^{p+q}$ is the entire style space. The bounded set of all styles provided in the industry at time $t$ is denoted as $\mathbb{Z}_{t}.$ It depends on the style of stores opened before time $t$ and their lifespan.
There are $m$ types of consumers who have different preference for the style. Their population structure is notated as $\mathbf{P}=\left(\mathbf{P}^{(1)},\mathbf{P}^{(2)},\ldots,\mathbf{P}^{(m)}\right).$ with $\mathbf{P}^{(j)}$ representing the proportion of the type $j$ consumers. Consumers will purchase products in different styles of stores with different probabilities. A consumer of type $j$th will assign a bounded continuous purchase probability density on $\mathbb{Z}_{t}$ at time $t$ and we denote this density function as $\rho_{jt}({\boldsymbol{z}}).$ We show in A that when the dimension of $\mathbb{Z}_{t}$ is smaller than the dimension of the entire style space $\mathbb{R}^{p+q}$, i.e. $\text{dim}\left(\mathbb{Z}_{t}\right)<p+q$, this probability density can still be defined, and we also show the definition of integrals related to it. Furthermore, we assume that styles with small difference may be indistinguishable. $\varepsilon$ represents the scale of indistinguishable difference. $\Gamma_{\varepsilon}(\mathbb{Z}_{t})$ is a $\varepsilon$-equipartition of $\mathbb{Z}_{t},$ which should satisfy that
@@
and $|\Delta|$ is constant for all $\Delta\in\Gamma_{\varepsilon}(\mathbb{Z}_{t}),$ where $\Phi$ is the empty set. Denote the subset in which style $\boldsymbol{z}$ is located as $\Delta_{t}(\boldsymbol{z},\varepsilon)\in\Gamma_{\varepsilon}(\mathbb{Z}_{t}).$ Given style ${\boldsymbol{z}}$ without analogs, a type $j$th consumer will purchase with a probability of $\tilde{\beta}_{jt}=\int_{\Delta_{t}(\boldsymbol{z},\varepsilon)}\rho_{jt}({\boldsymbol{z}_{1}})\mathrm{d}{\boldsymbol{z}_{1}}.$ Thus, the theoretical conversion rate for a store with style $\boldsymbol{z}$ at time $t$ can be calculated as
@@ 3.1 Style demand
First, we are going to discuss the utility function, the strategy, and the demand of consumers. We start by defining the scoring function of type $j$ consumers for the style vector. Similar to Berry and Haile (2014), we use a scoring function of the consumer type $j$ and remove the randomness, i.e. the $f$ function below is deterministic:
@@
The transform function maps elements in $\mathbb{U}$ to be positive and is strictly order-preserving. Some typical examples of transform functions are $e^{\kappa u}(\kappa>0)$ and $\ (C+C_{\varepsilon}-u)^{-\alpha}(\alpha>1,C_{\varepsilon}>0)$. (9) implies that the scoring function is a linear combination of style minus a penalty for deviation from tradition, transformulated and multiplied by the attenuation of price. The following proposition illustrates that each type of the consumers has an optimal preference style and that scoring decreases with distance from the optimal preference style.
@@
At time $t,$ there exists a unique optimal preference style for the $j$th type of consumer
@@
Under assumption 1, The $j$th type consumers’ optimization goal is
@@
where $K_{jt}$ is the level of purchase probability. In order to clarify some properties that will be mentioned later, we introduce several families of fake probability density of the form

### v2

@@ Abstract
The consumer store is ubiquitous and plays an important role in our everyday lives. It is an open question why stores usually have such short life cycles (typically around 3 years in China). This paper proposes a theoretical framework based on an equilibrium in style supply of stores and style demand of consumers to characterize store cash flow (revenue), leading to a strong explanation of this puzzle. In our model, we derive that the preference shifting of consumers is the main reason for the cash fl ow decreasing to its break-even line over time, while the visibility broadening leads to initial growth, resulting in a rainbow-shaped cash flow and its life cycle. Moreover, the intensified spatial competition will lead to an unexpected decrease in the store’s cash flow, or even closure. We calibrate our model with proprietary data of three Chinese stores from three representative industries and study the relationship between customers’ preference shifting and cash flow. To our knowledge, there have been no prior attempts to quantitatively model the life cycle of the store.
@@ 1 Introduction
In the vast forest of the marketplace, conventional listed companies are the towering trees, everlasting and long-living, while consumer stores (micro-enterprises) embody the plentiful small plants which naturally grow and die with the changing seasons. These stores, such as convenience stores, restaurants, and hotels, exist everywhere in our daily life, but little is known about them academically. Especially their frequent opening and closing, leaving a question of why these stores always have short lifespans. This paper develops a framework based on equilibrium in style to model stores’ cash flow over time, which can be calibrated by real data, and hence provides a capable explanation for this puzzle.
@@
The focus of examining a store’s life cycle is on the average level of cash flow and its changes over time [^1] In this paper, we do not distinguish the two words: the cash flow and the revenue.. Due to the necessary variable costs, such as costs of raw materials, rents, and salary, faced by a store, low levels of cash flow can lead to store closures. The generation of cash flow usually requires customers to pass through the front of the store, walk into the store and purchase goods or services (with a certain probability). Our model is built around this observation, taking into account several major factors that may affect a store’s cash flow. From a general perspective, we model the cash flow of one store via the following three steps, among which we propose novel concepts and assumptions to precisely characterize our framework.
@@
First, we decompose the cash flow of a store into the product of the average customer price (ACP), the conversion rate, and the flow of potential customers. Conversion rate is an important concept as it characterizes the probability of consumers consuming in a store. Perdikaki et al. (2012) define the conversion rate as the ratio of the number of transactions to that of the traffic and study the impact of traffic on cash flow. Furthermore, population structure is introduced to classify consumers into several types with different preference (such as based on their age, occupation, and income) with each type of consumers having its own conversion rate for a store. The overall conversion rate is the weighted average of conversion rates of each type of consumers.
@@
Second, we define style as the attributes of a store and its products, taking into account both the style demand of consumers and the style supply of store owners. On one hand, we capture the style preference of consumers by using the Cobb-Douglas (CD) utility function in Cobb and Douglas (1928) and the Berry-Levinsohn-Pakes (BLP) model in Berry et al. (1995). The latter model was the first to propose the utility of different types of consumers concerning the product style. The model of a time-varying form is later taken into account in Berry and Haile (2014). An important assumption is that consumers’ preferences shift over time. On the other hand, we explore the optimization goals of store owners to derive their optimal strategies. We assume that ACP and style of a given store is stable over time due to the price stickiness and the technical stickiness. But for some stores their styles can be updated over time. As a result, we find a Nash equilibrium, where we derive the mathematical form of the conversion rate for each type of consumers.
Third, we turn to the flow of potential customers. The modeling of it includes three points: the flow of foot traffic, distance decrease and visibility broadening. The foot traffic is assumed uniformly distributed around the store location with a constant density. Meanwhile, similar to Lucas and Rossi-Hansberg (2002), we assume a store’s attractiveness exponentially decreases with distance increasing. Furthermore, few consumers know about a store during its initial opening period and its visibility is broadening by a certain speed from the opening. Additionally, spatial competition is also considered. Intensive spatial competition leads to an unexpected drop in the flow of potential customers. To better illustrate, the debut of a new same-brand store near the initial store often prompts consumer defection. In cases where the local demand cannot sustain both establishments, the closure of one of the stores becomes inevitable.
@@
Finally, we obtain an equation of a store’s cash flow over time. In particular, we derive that when consumer preferences are not dispersed, the cash flow of a store over time is a rainbow-shaped curve. The increase in the cash flow curve can be interpreted as an increase in potential customers due to the visibility broadening, while the decrease can be interpreted as a decrease in the conversion rate due to consumers’ preferences shifting. We also use simulation method to illustrate some important implications. First, the simulation results indicate that (i) the visibility broadening speed mainly affects the store’s upfront cash flow; (ii) the decrease coefficient mainly affects the length of the store’s life cycle; and (iii) the foot traffic density mainly affects the macro-scale of cash flow over time. The results show that our model provides a strong explanation for the puzzle. As a conclusion, we provide a road map for our modeling framework in Figure 1.
@@
Cash flow $\text{CF}_{t}$ Average customer price $\theta_{t}$ Preferences flow of potential customers $N_{t}$ Product price $\boldsymbol{\theta}^{*}(\boldsymbol{z})$ Foot traffic density $u$ Attraction $\boldsymbol{q}_{t}$ Spatial competition Visibility broadening $k$ Attractiveness decrease $e^{-\delta\|\mathbf{x}\|}$ Consumer population structure $\mathbf{P}$ Score $F$, Utility $U$ Initial investment $I$ Store style: Product and storefront attributes $\mathbf{z}=(\mathbf{x},\boldsymbol{\xi})$ Equilibrium Purchase probability density $\rho^{*}$ Style set $\mathbb{Z}^{*}_{t}$ Consumer conversion rate $\tilde{\beta}_{t}$ Decrease $\nu$ Life cycle $T$
@@
In our empirical study, we use the real data from three Chinese stores from three industries in the real economy: the retail industry, the restaurant industry, and the service industry. We conduct a Non-linear Least Squares regression (NLS) of the adjusted cash flow on the time scale to calibrate the visibility broadening speed, the initial conversion rate, and the decrease coefficient. The regression results uncover that the estimates are all statistically significant at a 1% confidence level. Finally, we plot the real data and the fitted data to show that our model well captures the trend of the store’s cash flow, which strongly justifies the explanation of our model.
Our model incorporates several key ingredients of interest in literature, including product specialization, customers’ preferences, and purchase conversion rate. Burdett and Judd (1993) establish the equilibrium for dispersed prices under certain cases. Menzio (2023) discusses the relationship between product specialization to customers’ preferences, search friction decrease, and price dispersion. Menzio (2023) finds that the increase in product specialization helps sellers cater to customers’ preference shifting, which leads to a higher store profit and also a higher customer surplus. The search friction can be seen as the visibility broadening in our paper. Kashyap (1995) uses real data to illustrate the effects of the size, frequency, and synchronization of price changes for twelve selected retail goods. In the field of marketing science, Lam et al. (2001) propose a framework to carefully classify the conversion rate into three effects of attraction (store-entry), conversion (purchase decision), and spending (purchase amount). They formulate a joint model of four simultaneous equations using front traffic, store traffic, number of store transactions, and store sales as the endogenous variables. Zimmermann and Auinger (2023) study a digital retailer and adopt a conversion rate optimization framework to increase sales.
The paper is organized as follows. Section 2 proposes the model environment of our paper. The demand of consumers and the supply of stores are well defined in Section 3. Section 4 defines the concept of preference shifting and derives a Nash equilibrium. The spatial competition is described and we close the model in Section 5. The full model is presented in Section 6. We provide empirical evidence in Section 7. Section 8 provides further discussions. Section 9 concludes.
@@ 2 Model Environments
Store. We study the typical consumer store in this paper. A (consumer) store is a type of micro-enterprise, usually possessing the following features: (i) Initial monetary investment is small. For example, it costs the owner around 400,000 Chinese yuan to invest in a luckin coffee store in Beijing, China. (ii) Routine expenditures include rent and salary, which form a surprisingly stable ratio to daily income. For instance, the ratio of daily rent to daily cash flow is around 6.17% for a typical Chinese convenience store in 2022, according to the annual report of KPMG and China Chain-Store & Franchise Association (CCFA). (iii) The store has a high-frequency cash flow. For example, a Family Mart store typically has 1,000 orders (purchases) per day. (iv) The store caters to people’s daily demands, which encompasses the restaurant, retail, and service industries. (v) Most consumers of stores purchase goods and services in-store, especially the stores in the service industry. Therefore, the location of the store impacts their financial performance. (vi) There exists perfect competition (over 100 million stores in China) in each industry. In short, our focus is not on the conventional company, but on the store, which is usually the following three types: direct-sale store, franchised outlet, and ’mom-and-pop store’.
Cash flow decomposition. In this paper, we fix one store on time set $\mathcal{T}.$ The cash flow CF is a non-negative function on $\mathcal{T}.$ Then, we model the cash flow at time $t$ by a product of the average customer price (ACP) $\theta_{t}$, the flow of potential customers $N_{t}$ and the conversion rate $\beta_{t}$, which can be formulated as
@@
Attributes and Styles. Consider an economy corresponding to a consumer industry on time set $\mathcal{T}=\mathbb{R}.$ At any time $t\in\mathcal{T}$, there are new stores opened with different initial investments. Let the set of all initial investments be $\mathcal{I}\subseteq\mathbb{R}^{+},$ which is a bounded set that does not change over time. Therefore, stores are infinitely numerous in this economy. The style of a store is the combination of its products and storefront. We assume that a store provides one representative product (commodity), characterized by a constant price $\theta$ and product attributes $\boldsymbol{x}\in\mathbb{R}^{p}$, which reflect the features of product size, taste, shape, etc. Similar assumption is mentioned or implied in Chamberlin (1933), Dixit and Stiglitz (1977) and Romer (1986). Moreover, we denote the storefront attributes by $\boldsymbol{\xi}\in\mathbb{R}^{q}$, which reflect the implicit influencing factors of store service, decoration, reputation, etc. We define the style vector for a store $\boldsymbol{z}=(\boldsymbol{x}^{\prime},\boldsymbol{\xi}^{\prime})^{\prime},$ and $\mathbb{R}^{p+q}$ is the entire style space. The bounded set of all styles provided in the industry at time $t$ is denoted as $\mathbb{Z}_{t}.$ It depends on the styles of stores opened before time $t$ and their lifespans.
There are $m$ types of consumers who have different preferences for style. Their population structure is notated as $\mathbf{P}=\left(\mathbf{P}^{(1)},\mathbf{P}^{(2)},\ldots,\mathbf{P}^{(m)}\right)$ with $\mathbf{P}^{(j)}$ representing the proportion of the $j$th type of consumers. Consumers will purchase products in different styles of stores with different probabilities. A consumer of type $j$th will assign a bounded continuous purchase probability density on $\mathbb{Z}_{t}$ at time $t$ and we denote this density function as $\rho_{jt}({\boldsymbol{z}}).$ We show in A the details about the definition of $\rho_{jt}$ and the integrals with respect to it. Furthermore, we assume that styles with small differences may be indistinguishable. $\varepsilon$ represents the scale of indistinguishable difference. $\Gamma_{\varepsilon}(\mathbb{Z}_{t})$ is a $\varepsilon$-equipartition of $\mathbb{Z}_{t},$ which should satisfy that
@@
and $|\Delta|$ is constant for all $\Delta\in\Gamma_{\varepsilon}(\mathbb{Z}_{t}),$ where $\Phi$ is the empty set. Denote the subset in which style $\boldsymbol{z}$ is located as $\Delta_{t}(\boldsymbol{z},\varepsilon)\in\Gamma_{\varepsilon}(\mathbb{Z}_{t}).$ Given style ${\boldsymbol{z}}$ without analogs, a $j$th type consumer will purchase with a probability of $\tilde{\beta}_{jt}=\int_{\Delta_{t}(\boldsymbol{z},\varepsilon)}\rho_{jt}({\boldsymbol{z}_{1}})\mathrm{d}{\boldsymbol{z}_{1}}.$ Thus, the theoretical conversion rate for a store with style $\boldsymbol{z}$ at time $t$ can be calculated as
@@ 3.1 Style demand
First, we are going to discuss the utility function, the strategy, and the demand of consumers. We start by defining the scoring function of type $j$ consumers for style. Similar to Berry and Haile (2014), we use a function of consumer type, time, and attributes but remove the randomness, i.e. the $f$ function below is deterministic:
@@
The transform function maps elements in $\mathbb{U}$ to be positive and is strictly order-preserving. Some typical examples of transform functions are $e^{\kappa u}(\kappa>0)$ and $\ (C+C_{\varepsilon}-u)^{-\alpha}(\alpha>1,C_{\varepsilon}>0)$. (9) implies that the scoring function is a linear combination of style minus a penalty for deviation from tradition, transformed and multiplied by the attenuation of price. The following proposition illustrates that each type of consumers has an optimal preference style and that scoring decreases with distance from the optimal preference style.
@@
At time $t,$ there exists a unique optimal preference style for the $j$-th type of consumer
@@
Under Assumption 1, The $j$-th type consumers’ optimization goal is
@@
where $K_{jt}$ is the level of purchase probability. In order to clarify some properties that will be mentioned later, we introduce several families of the generalized probability density of the form